# Volumetric Surfaces: Representing Fuzzy Geometries with Multiple Meshes

### v1

@@ 4.2 Surface Blending
Transparency Attenuation: Blending different hard surfaces might result in clear cut-offs at their borders, especially noticeable from test views. We prefer smoother transitions towards full transparency. We achieve this by multiplying predicted transparencies with a weight $\alpha_{\text{w}}$ that depends on the angle between the view direction and the surface normal, downweighting the contribution from grazing angles. Specifically, we use
$\alpha_{\text{w}}=2\cdot\text{Sigmoid}(10\cdot|\mathbf{v}\cdot\mathbf{n}|)-1\text{.}$ (8)
@@ 4.3 Training and Baking
Our method is composed of two main stages. First, we optimize an implicit representation of $k$ surfaces and their appearances. Then, we fine-tune resolution-bounded neural textures over their explicit (meshes) representation. In the following, we describe both phases in detail.
4.4 Implicit Surfaces
We first train a standard NeuS-like [45] model for 100,000 iterations, during which $\beta$ is scheduled to go from large $\phi_{\beta_{1}}$ to thin densities $\phi_{\beta_{2}}$ exponentially. In practice, we find that training the $k$-SDF model from scratch with predicted transparencies leads to the presence of fully transparent geometry in the reconstruction, which we want to avoid. The reconstructed main surface acts as an anchor for the initialization of the remaining $k-1$ support surfaces, represented as shells from the main surface uniformly distanced by $\Delta o$. The mathematical formulation of $k$-SDF (Section 4.1) allows us to initialize support surfaces on the inside and on the outside of the main SDF; however, we observed that initializing them all on the inside leads to more stable runs. This is because the reconstructed main surface tends to be conservative, acting as an outer shell of the scene content. In practice, this leads support surfaces to be initialized on the outside to be modeled as fully transparent in the very first iterations of joint training. As transparent surfaces do not contribute to the final image, gradient flow is inhibited and the unused support surfaces do not recover. While we have observed the optimization to be robust under different values of $\Delta o$, we found a good balance by setting it by the logistic distribution function $\phi_{\beta}$ standard deviation $(1/\beta)\pi/\sqrt{3}$. The full model, composed of all surfaces, is trained for 50,000 iterations starting from $\phi_{\beta_{2}}$, until a $\phi_{\beta_{3}}$ value for which all surfaces are modeled as peaked densities. When this happens, our rendering process collapses to a simple blending of hard surfaces (Figure 2(c)), making the reconstructed set of implicit surfaces optimal for the later steps.
During training, we enforce two additional losses on all surfaces. The Eikonal loss [13], calculated on points in the vicinity of the zero-level sets and on randomly sampled points, and a curvature loss [35] on near-surfaces points, to push the optimization toward smooth solutions. We minimize $\mathcal{L}=\mathcal{L}_{\text{c}}+\lambda_{\text{e}}\mathcal{L}_{\text{e}}+\lambda_{\text{s}}\mathcal{L}_{\text{s}}$, where $\mathcal{L}_{\text{c}}$ is the standard $l_{1}$ pixel-wise color loss, $\mathcal{L}_{\text{e}}$ is the Eikonal loss, and $\mathcal{L}_{\text{s}}$ is the curvature term. We use weights $\lambda_{\text{e}}=0.04$ and $\lambda_{\text{s}}=0.65$.
@@ 4.4 Implicit Surfaces
Occupancy grid: As training progresses and densities peak, our volumetric rendering samples are gradually positioned closer to the zero-level sets of the signed distance functions, since points farther away would be in empty space. To do so, we compute per-surface binary occupancy values by describing as occupied voxels those whose center point’s $\left|d\right|$ value, together with the current $\beta$ and the voxel’s space diagonal, would allow any point in the voxel volume to have density above a $10^{-4}$ threshold. Per-surface occupancy values are then aggregated with an or operation, resulting in a single binary occupancy grid. Rays traverse the grid to sample $n$ points uniformly in occupied space. In our experiments, we used a grid of resolution $256^{3}$.
@@
Importance sampling: We adopt hierarchical sampling from NeuS [45], extending it to the $k$-surfaces case. Starting from the $n$ uniform samples, each SDF is evaluated to compute $k$ CDFs; these are summed together and normalized. The resulting probability distribution is used to sample $m/2$ additional points that are added to the previous $n$. The whole operation is then repeated on the expanded set of points to add $m/2$ more samples. The resulting number of samples per ray is then $n+m$. The kernel size $\beta$ used to compute weights at the second iteration is the same as used during rendering, $\beta$ at the second iteration is half the value.
@@
Mesh baking: We extract $k$-SDF zero-level sets as high-resolution meshes using marching cubes [21]. and heavily simplify them [12] to 0.02% of the original number of triangles (approximately 2 MB each for synthetic scenes), meeting the strict computing and memory requirements of mobile hardware. Finally, we compute UV atlases for these lightweight meshes with xatlas [52]. This is a fundamental step for the next stage, in which we train per-mesh neural SH textures, which are later baked as explicit textures.
@@
Figure 5: Bilinear interpolation in our mixed-resolution textures. Instead of directly querying the blue 2D point, we predict the values of its surrounding texel centers (red points) and bilinearly interpolate them. This makes the neural texture texture-dependent, with predefined width $W$ and height $H$.
@@
Table 1: Results are averaged across all testing scenes. Highlighted baselines meet the compute and/or memory requirements to run on general-purpose hardware (via WebGL) without modifications, as discussed in Section 2.1. Methods marked with a $\star$” show results from original papers, as code is unavailable. MobileNeRF and Instant-NGP results on Shelly from AdaptiveShells. DTU is evaluated on masked foreground. PermutoSDF trained until densities are fully peaked ($\phi_{\beta_{3}})$. Metrics not provided by a baseline are denoted with “—”. Please note that the Shelly dataset, as released by the authors, has a large overlap of views between test and training sets. All our experiments were conducted on a cleaned version of the dataset, free of this problem. However, a fair comparison with baselines whose code is not public remains difficult. Our per-scene metrics are reported in the supplementary material.
Shelly [47] Custom (ours + [38]) DTU [16] (scans 83, 105)
@@
3DGS [17] 35.443 0.964 0.064 37.345 0.982 0.147 38.060 0.989 0.086
Instant-NGP [26] 33.220 0.922 0.125 31.133 0.935 0.132 38.245
PermutoSDF [35] 29.850 0.950 0.129 33.306 0.961 0.193 36.312 0.988 0.098
AdaptiveShells [47] 36.020 0.954 0.079
QuadFields [38] 35.130 0.954 0.073
MobileNeRF [5] 31.620 0.911 0.129
3-Mesh 33.247 0.977 0.118 34.843 0.970 0.175 36.242 0.984 0.095
5-Mesh 34.098 0.979 0.114 35.419 0.974 0.171 36.504 0.985 0.089
7-Mesh 34.331 0.980 0.113 35.605 0.976 0.169 36.624 0.987 0.087
9-Mesh 34.220 0.981 0.113 35.716 0.978 0.167 36.826 0.986 0.086
4.5 Mesh Texturing
Lastly, we train our final per-surface appearance model based on neural textures. A neural texture is implemented as a hash-grid with input dimension 2 and a small decoder MLP with output dimension dependent on the SH degree it models. The $i$-th neural texture predicts $\overline{\mathbf{sh}_{i}}\in\mathbb{R}^{4\times c_{\text{deg}}}$, with $c=\{1,3,5,7\}$. The outputs of all neural textures are concatenated over the coefficient dimension ($\mathbf{sh}\!\in\!\mathbb{R}^{4\times(\text{deg}+1)^{2}}$) and decoded with view direction $\mathbf{v}$ to predict RGBA. Specifically, we discretize our neural textures by grounding them to a fixed target baking resolution. In this setting, each point $\mathbf{x}\!\in\!\mathbb{R}^{3}$ on the surface of a mesh is mapped to a point $\overline{\mathbf{x}}\!\in\![0,1]^{2}$ in UV space. SH coefficients at $\overline{\mathbf{x}}$ are predicted as the result of a bilinear interpolation of predictions at neighboring texels centers $\{\overline{\mathbf{x}}_{i}\}_{1,\ldots,4}$, as visualized in Figure 5. This effectively mimics how OpenGL fragment shaders access and interpolate texture values; by doing so, we optimize our texture values such that when displayed in our real-time renderer, images match – up to numerical precision – those synthesized by our differentiable renderer. This final phase is trained for 15,000 iterations without additional losses.
Mixed resolutions: Storing all textures at full resolution ($2048^{2}$) is impractical on mobile devices, as it would require approximately 0.5 GB per mesh. We propose scaling texture resolution proportionally to the spherical harmonics (SH) coefficient degree: the base color is modeled at the highest resolution ($2048^{2}$), while the highest SH degree coefficients are modeled at the lowest resolution ($256^{2}$). This approach significantly reduces the memory footprint to approximately 14 MB per mesh without compromising image quality, as shown in Table 5.
Squeezing and quantization: Neural textures predicted values $\mathbf{sh}$ are continuous and defined in an unbounded range. Before storing them, they must be squeezed to the unit range via $\hat{\mathbf{sh}}=\text{Sigmoid}(\mathbf{sh})$. Training must account for quantization to the discrete $[0,255]$ range when storing textures as images. Following MERF [33], we apply the function $q(x)=\text{round}(255x)/255$ to the predicted and squeezed coefficients. Before rendering, we re-scale value to a hyper-parameters controlled range $[v_{\text{min}},v_{\text{max}}]$:
$\mathbf{sh}=v_{\text{min}}+(v_{\text{max}}-v_{\text{min}})q(\hat{\mathbf{sh}})$ (9)
In practice, we observed a range of $[-15,15]$ to work well on all scenes.
Texture baking: Our resolution-grounded neural textures are easy to bake, as this only requires predicting values at texel centers and storing them. Baking results in $(\text{deg}+1)^{2}$ PNG images, where the $i$-th image stores the $i$-th coefficient of RGBA channels. The fully baked representation can finally be visualized in our WebGL renderer, which rasterizes all meshes in a fixed order and blends them in the final frame buffer before displaying the result on screen.
5 Evaluation
Our most important baselines are methods targeting real-time rendering. We identified 3DGS and MobileNeRF as our main competitors; 3DGS represents the fastest volumetric representation, while MobileNeRF is a surface-based method pioneering mobile neural graphics rendering. We compare them in terms of quality (Table 1), speed and memory footprint (Table 2). Additionally, we compare with other baselines not designed for rendering on general-purpose hardware to provide a more comprehensive overview of current methods. Aiming to represent fuzzy geometries with high quality, we focus our evaluation on object-centric datasets featuring fuzzy structures. We present results from synthetic benchmark datasets like Shelly [47], plush objects from real-world tabletop scenes (DTU [16]), and additional synthetic custom scenes (our plushy and hairy_monkey from [38]). Our method consistently achieves higher image quality than surface-based competitors and renders faster than 3DGS, even with the maximum number of surfaces. Using seven layers offers a good balance between image quality, model size, and rendering speed, as quality tends to degrade with nine meshes under the same number of training iterations. This occurs because deeper surfaces contribute less to the final pixel color than outer surfaces, leading to scaled gradient magnitudes and making them harder to optimize. Figure 6 pictures the trade-offs between frame rate and model size. While our method falls short of the latest volume-based baselines in image quality, our baked representation provides a favorable balance between quality and speed, rendering much faster on non-specialized hardware. We refer to our supplementary material for additional results visualizations.
@@ 5 Evaluation
Figure 6: Framerate vs quality comparison (laptop results from Table 2). The radius of each circle represents the memory footprint of each model as stored on disk.
Table 2: 3DGS with spherical harmonics up to degree 2, ours up to degree 3. Metrics averaged over scenes of the Shelly [47] dataset. Framerate is measured on close-up views at HD (720p) resolution on low-power laptop (Dell XPS 13 i5) (marked with $\star$) and smartphone (Samsung A52s) (marked with $\diamond$), on respective WebGL renderers (when available); memory footprint as stored on disk.
Method FPS ${}^{\diamond}\uparrow$ FPS ${}^{\star}\uparrow$ MB $\downarrow$
3DGS [17] 8.0 18 53.6
QuadFields [38] 1213.0
Our 3-Mesh 55.0 152 43.8
Our 5-Mesh 40.0 97 69.1
Our 7-Mesh 27.5 72 97.6
Our 9-Mesh 21.5 51 133.6
@@
Figure 7: Qualitative comparisons between b) PermutoSDF [35], a state-of-the-art implicit surface-based method, and c) our Volumetric Surfaces demonstrate that our approach convincingly represents fuzzy objects. This is achieved by trading a) high-frequency geometry for the number of integration points, which are found by rasterizing smooth, lightweight meshes defined as d) shells around the object and traversed in a fixed order. Volumetric Surfaces enable fast rendering of fuzzy geometries on general-purpose hardware with image quality approaching the latest volumetric representations (Table 1). Scenes from Shelly [47] and QuadFields [38].

### v2

@@ 4.2 Surfaces Blending
Transparency Attenuation: Blending different hard surfaces might result in clear cut-offs at their borders, especially noticeable from test views. We prefer smoother transitions towards full transparency. We achieve this by multiplying predicted transparencies with a weight $\alpha_{\text{w}}$ that depends on the angle between the view direction and the surface normal, downweighting the contribution from grazing angles. Specifically, we use: $\alpha_{\text{w}}=2\cdot\text{Sigmoid}(10\cdot|\mathbf{v}\cdot\mathbf{n}|)-1$.
$a$) Refer to caption
$b$) Refer to caption
[…]
@@ 4.3 Training and Baking
Our method is composed of two main stages. First, we optimize an implicit representation of $k$ surfaces and their appearances. Then, we fine-tune resolution-bounded neural textures over their explicit (meshes) representation. Our hyperparameters are robust, ensuring consistent performance across all scenes and datasets. In the following, we describe both phases in detail.
Implicit Surfaces: We begin by training a standard NeuS-like [54] model for 100k iterations, during which $\beta$ is exponentially scheduled from large densities $\phi_{\beta_{1}}$ to thinner densities $\phi_{\beta_{2}}$. In practice, training the $k$-SDF model from scratch with predicted transparencies often introduces fully transparent additional geometry in the reconstruction. This opaque pre-training step helps prevent such artifacts. The reconstructed surface serves as an anchor for initializing the remaining $k\!-\!1$ support surfaces, which are represented as shells uniformly spaced from the main surface by $\Delta o$. The mathematical formulation of $k$-SDF (Section 4.1) allows for initializing support surfaces both inside and outside the main SDF. However, we find that initializing all support surfaces on the inside increases model capacity, leading to better results (Section 5.1).
While we have observed the optimization to be robust under different values of $\Delta o$, we found a good balance by setting it by the logistic distribution function $\phi_{\beta_{2}}$ standard deviation: $\Delta o=(1/\beta_{2})\pi/\sqrt{3}$. This ensures that surface densities only partially overlap (e.g., Figure 4). The $k$-SDF model is trained for 50k iterations starting from $\phi_{\beta_{2}}$, until a $\phi_{\beta_{3}}$ value for which all surfaces are modeled as peaked densities. When this happens, our rendering process collapses to a simple blending of hard surfaces (Figure 1c), making the reconstructed set of implicit surfaces optimal for the later steps. During both training phases, we apply two additional losses to all surfaces. The Eikonal loss $\mathcal{L}_{\text{e}}$ [16], calculated on points in the vicinity of the zero-level sets and on randomly sampled points, and a curvature loss $\mathcal{L}_{\text{s}}$ [44] on near-surfaces points, to push the optimization toward smooth solutions. We minimize $\mathcal{L}=\mathcal{L}_{\text{c}}+\lambda_{\text{e}}\mathcal{L}_{\text{e}}+\lambda_{\text{s}}\mathcal{L}_{\text{s}}$, where $\mathcal{L}_{\text{c}}$ is the standard $L_{1}$ pixel-wise color loss, $\lambda_{\text{e}}=0.04$ and $\lambda_{\text{s}}=0.65$.
@@
Occupancy Grid: As implicit surfaces training progresses and densities peak, our volumetric rendering samples are gradually positioned closer to the zero-level sets of the signed distance functions, since points farther away would be in empty space. To do so, we compute per-surface binary occupancy values by describing as occupied voxels whose center point’s $\left|d\right|$ value, together with the current $\beta$ and the voxel’s space diagonal, would allow any point in the voxel volume to have density above a $10^{-4}$ threshold. Per-surface occupancy values are then aggregated with an or operation, resulting in a single binary occupancy grid. Rays traverse the grid to sample $n$ points uniformly in occupied space. In our experiments, we used a grid of resolution $256^{3}$.
@@
Importance Sampling: During implicit surfaces training, we adopt hierarchical sampling from NeuS [54], extending it to the $k$-surfaces case. Starting from the $n$ uniform samples, each SDF is evaluated to compute $k$ CDFs; these are summed together and normalized. The resulting probability distribution is used to sample $m/2$ additional points that are added to the previous $n$. The whole operation is then repeated on the expanded set of points to add $m/2$ more samples. The resulting number of samples per ray is then $n+m$. In the first iteration, the kernel size $\beta$ is half of that used during rendering, while in the second, it matches it.
@@
Baking Meshes: After the implicit surfaces optimization, we extract $k$-SDF zero-level sets as high-resolution meshes using marching cubes [30] and simplify them significantly [15, 31] to 0.02% of the original triangle count (approximately 2 MB per mesh for synthetic scenes), meeting the strict computing and memory constraints of mobile hardware. Finally, we generate UV atlases for these lightweight meshes using xatlas [61]. This step is essential for training per-mesh neural SH textures, which are later baked into explicit textures.
@@
Figure 5: Bilinear interpolation in our mixed-resolution textures. Instead of querying the blue 2D point directly, we predict the values at its surrounding texel centers (red points) and bilinearly interpolate them. This anchors the neural texture to a predefined target resolution ($W$, $H$).
Texturing Meshes: Finally, we train a per-surface view-dependent appearance model for color and transparency using neural textures. A neural texture is implemented as a hash-grid with input dimension 2 and a small decoder MLP with output dimension dependent on the SH degree it models. The $i$-th neural texture predicts $\overline{\mathbf{sh}_{i}}\in\mathbb{R}^{4\times c_{\text{deg}}}$, with $c=\{1,3,5,7\}$. The outputs of all neural textures are concatenated over the coefficient dimension ($\mathbf{sh}\!\in\!\mathbb{R}^{4\times(\text{deg}+1)^{2}}$) and decoded with view direction $\mathbf{v}$ to predict RGBA. Specifically, we discretize our neural textures by grounding them to a fixed target baking resolution. In this setting, each point $\mathbf{x}\!\in\!\mathbb{R}^{3}$ on the surface of a mesh is mapped to a point $\overline{\mathbf{x}}\!\in\![0,1]^{2}$ in UV space. SH coefficients at $\overline{\mathbf{x}}$ are predicted as the result of a bilinear interpolation of predictions at neighboring texel centers $\{\overline{\mathbf{x}}_{i}\}_{1,\ldots,4}$, as visualized in Figure 5. This effectively mimics how OpenGL fragment shaders access and interpolate texture values; by doing so, we optimize our texture values such that when displayed in our real-time renderer, images match – up to numerical precision – those synthesized by our differentiable renderer. This final phase is trained for 15k iterations using the $L_{1}$ pixel-wise color loss.
Table 1: Sweep over the number of meshes and target neural textures resolution. Mixed resolutions are described in Figure 5. Increasing neural texture resolution improves reconstruction quality up to $1024^{2}$, but performance declines beyond that. Modeling all textures at the maximum resolution ($2048^{2}$) is counterproductive, as results indicate that a mixed-resolution approach yields better image quality while reducing memory usage. Results on the khady scene from Shelly [56]; we highlight the best, second best and third best.
PSNR $\uparrow$
Method mixed 20482 10242 5122 2562
5-Mesh 29.88 29.88 29.91 29.82 29.54
7-Mesh 29.97 29.97 30.02 29.94 29.65
9-Mesh 29.96 29.93 29.97 28.91 29.62
Mixed Resolutions: Storing all textures at full resolution ($2048^{2}$) is impractical, as it would require approximately 0.5 GB per mesh. We propose scaling texture resolution proportionally to the spherical harmonics (SH) coefficient degree: the base color is modeled at the highest resolution ($2048^{2}$), while the highest SH degree coefficients are modeled at the lowest resolution ($256^{2}$). This approach significantly reduces the memory footprint to approximately 14 MB per mesh without compromising image quality (Table 1).
Squeezing and Quantization: Neural textures predicted values are continuous and unbounded. Before storing them, they must be squeezed to the unit range via $\hat{\mathbf{sh}}=\text{Sigmoid}(\mathbf{sh})$. Training must account for quantization to the discrete $[0,255]$ range when storing textures as images. Following MERF [42], we apply the function $q(x)=\text{round}(255x)/255$ to the predicted and squeezed coefficients. Before rendering, we re-scale value to a hyper-parameters controlled range $[v_{\text{min}},v_{\text{max}}]$: $\mathbf{sh}=v_{\text{min}}+(v_{\text{max}}-v_{\text{min}})q(\hat{\mathbf{sh}})$. We observed a range of $[-15,15]$ to work well on all scenes.
Baking Textures: Lastly, we bake our resolution-grounded neural textures, which is straightforward as it only requires predicting values at texel centers and storing them. Baking results in $(\text{deg}+1)^{2}$ PNG images, where the $i$-th image stores the $i$-th coefficient of RGBA channels. The fully baked representation can finally be visualized in our WebGL renderer, which rasterizes all meshes in a fixed order and blends them in the final frame buffer before displaying it on screen.
5 Evaluation
Table 2: Framerate is measured on close-up views at HD (720p) resolution on low-power smartphone (Samsung A52s) (marked with $\diamond$) and laptop (Dell XPS 13 i5) (marked with $\star$), on respective WebGL renderers; memory footprint as stored on disk. 3DGS [24] with spherical harmonics of degree 2, ours of degree 3. Metrics averaged over scenes of the Shelly [56] dataset. We present further qualitative comparisons in the supplementary material.
Method FPS ${}^{\diamond}\uparrow$ FPS ${}^{\star}\uparrow$ PSNR $\uparrow$ MB $\downarrow$
MobileNeRF [7] 24 35 29.30 194
3DGS-50K 20 160 32.73 12
3DGS-75K 13 115 33.05 18
3DGS [24] 8 18 35.44 57
3-Mesh 65 145 33.39 46
5-Mesh 55 90 34.25 77
7-Mesh 42 70 34.50 110
9-Mesh 35 55 34.38 140
@@ 5 Evaluation
Table 3: Results are averaged across all testing scenes; we highlight the best, second best and third best results. Highlighted baselines meet the compute and/or memory requirements to run on general-purpose hardware (via WebGL) without modifications, as discussed in Section 2. QuadFields [47] fails to meet real-time requirements due to its reliance on specialized ray-tracing acceleration hardware and its high memory consumption (e.g., Shelly [56] scenes require an average of 1213 MB). Our results are directly computed from our WebGL real-time render on our final baked representation. Methods marked with a $\star$” show results from original papers, as code is unavailable at the time of writing. Instant-NGP [35] results on Shelly from Wang et al. [56]. DTU [21] is evaluated on masked foreground. PermutoSDF [44] trained until densities are fully peaked ($\phi_{\beta_{3}})$. Metrics not provided by a baseline are denoted with “—”. Please note that the Shelly dataset, as released by the authors, has a large overlap of views between test and training sets. All our experiments were conducted on a cleaned version of the dataset, free of this problem. However, a fair comparison with baselines whose code is not public remains difficult. Our per-scene metrics are reported in the supplementary material.
Shelly [56] Custom (plushy + [47]) DTU [21] (scans 83, 105)
@@
3DGS [24] 35.44 0.975 0.089 37.34 0.982 0.147 38.06 0.989 0.086
Instant-NGP [35] 33.22 0.922 0.125 31.13 0.935 0.132 38.24
PermutoSDF [44] 29.85 0.950 0.129 33.31 0.961 0.193 36.31 0.988 0.098
AdaptiveShells [56] 36.02 0.954 0.079
QuadFields [47] 35.13 0.954 0.073
MobileNeRF [7] 29.30 0.939 0.150 30.89 0.942 0.195
3-Mesh 33.39 0.978 0.115 35.00 0.970 0.171 36.41 0.985 0.092
5-Mesh 34.25 0.980 0.110 35.45 0.975 0.172 36.87 0.986 0.085
7-Mesh 34.50 0.981 0.109 35.63 0.977 0.169 36.77 0.987 0.084
9-Mesh 34.38 0.981 0.110 35.74 0.978 0.167 37.17 0.987 0.083
@@
Figure 6: Frame rate vs. image quality comparison (smartphone results $\diamond$ from Table 2). The radius of each circle represents the memory footprint as stored on disk. The vertical dashed line marks the required frame rate for real-time rendering (30 FPS).
@@
Figure 7: Qualitative comparisons between ($b$) PermutoSDF [44], a state-of-the-art implicit surface-based method, and ($c$) our Volumetric Surfaces demonstrate that our approach convincingly represents fuzzy objects. This is achieved by trading ($a$) high-frequency geometry for the number of integration points, which are found by rasterizing smooth, lightweight meshes defined as ($d$) shells around the object and traversed in a fixed order. Volumetric Surfaces enable fast rendering of fuzzy geometries on general-purpose hardware with image quality approaching the latest volumetric representations (Table 3). Scenes from Shelly [56] and QuadFields [47].
Our key baselines focus on real-time rendering, with 3DGS and MobileNeRF as primary competitors. 3DGS represents the fastest volumetric approach, while MobileNeRF pioneers surface-based neural graphics for mobile devices. We compare to these methods in terms of quality, speed, and memory footprint on a low-power laptop (Dell XPS 13 i5) and a smartphone (Samsung A52s) (Table 2). Additionally, we compare with other baselines not designed for general-purpose hardware rendering to provide a broader overview of current methods (Table 3). Aiming at high-quality representation of fuzzy geometries, we focus our evaluation on object-centric datasets with prominent fuzzy structures. We present results from synthetic benchmark datasets like Shelly [56], plush objects from real-world tabletop scenes (DTU [21]), and additional synthetic custom scenes (our plushy and hairy monkey from Sharma et al. [47]).
Our method, tested on renderings from our real-time WebGL renderer, consistently delivers higher image quality than surface-based competitors and renders faster than 3DGS. We observe that our adaptive shell spacing clusters surfaces around solid structures while maintaining greater separation in volumetric regions (Figure 7). Using seven layers offers a good balance between image quality, model size, and rendering speed, as quality tends to degrade with nine meshes under the same number of training iterations. This happens as deeper surfaces contribute less to pixel color, reducing gradient magnitudes and slowing optimization. Figure 6 illustrates the trade-offs between frame rate on a low-cost smartphone and model size. Notably, 3DGS fails to meet real-time requirements even when the number of Gaussians is capped during optimization, which leads to a substantial loss in quality.
For comparability, the frame rate of 3DGS is measured using a widely adopted web viewer [22]. This implementation skips per-frame sorting to prevent slowdowns, using occasional CPU sorting instead. This results in noticeable popping artifacts during rapid camera rotation and suboptimal performance on mobile devices. Our sorting-free method avoids these issues. While our method falls short of the latest volume-based baselines in image quality, our baked representation provides a favorable balance between quality and speed, rendering much faster on non-specialized hardware. We refer to our supplementary material for additional results visualizations.
Table 4: Ablation studies over intermediate results (5-SDF, implicit representation, before meshes baking) and on final results (5-Mesh, baked, real-time rendering assets). See Section 5.1 for explanations. 5-Mesh achieves higher quality than the 5-SDF due to its fixed geometry and surface-constrained appearance model; the SDF representation suffers from the stochastic nature of ray sampling, forcing the appearance model to allocate capacity to off-surface elements. Results averaged over Shelly [56].
5-SDF 5-Mesh
Ablation PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Full 32.05 0.964 0.130 34.25 0.980 0.110
1) w/o view-dep. $\alpha$ 31.75 0.962 0.131 32.71 0.975 0.120
2) w/o curvature $\mathcal{L}_{\text{s}}$ 33.02 0.971 0.117 33.41 0.980 0.110
3) w/o $\alpha_{\text{w}}$ 32.11 0.966 0.130 33.96 0.980 0.113
4) w. const. $\Delta o$ 30.09 0.950 0.129
5) w. outer init. 30.85 0.955 0.121